# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v1

@@ 3.2 Existence of optimal relaxed controls
For $N\geq 1$, define $U^{N}(t,\omega):=\{u\in U(t,\omega);~{}|u|\leq N\}$. Then, for each $(t,\omega)\in[0,T]\times\Omega$, $U^{N}(t,\omega)$ increases to $U(t,\omega)$. We also introduce $B^{N}=\{(t,\omega)\in[0,T]\times\Omega;~{}U^{N}(t,\omega)\neq\varnothing\}$. Since $U(t,\omega)$ is bounded and closed, consequently, it is also a compact subset of $U\subset\mathbb{R}^{l}$. Then, there exists a measurable selection $\alpha^{N}:B^{N}\to U$ such that $\alpha^{N}_{t}(\omega)\in U^{N}(t,\omega)$ for $(t,\omega)\in[0,T]\times\Omega$ and $\alpha^{N}$ is $\mathcal{B}([0,t])\otimes\mathcal{F}_{t}/\mathcal{B}(\mathbb{R}^{l})$ measurable due to Leese [28]. Hence, we can define an $\mathbb{F}$-progressively measurable process $\alpha=(\alpha_{t})_{t\in[0,T]}$ by $\alpha_{t}(\omega)=\sum_{N=1}^{\infty}\alpha^{N}_{t}(\omega)\mathbf{1}_{B_{N}\backslash B_{N-1}}(t,\omega)$.
@@
+$\infty$+ ∞ $>\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},\mathcal{L}(X_{t}),u)q_{t}(\mathrm{d}u)\mathrm{d}t\right]\geq\mathbb{E}^{R}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}(X_{t}),\alpha_{t})\mathrm{d}t\right]$
@@
This, together with Lemma 3.2, indicates that $\mathbb{E}^{R}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. On the other hand, using the construction of $\alpha$ again, we deduce that, for every $g\in C_{b}^{2}(\mathbb{R}^{n})$ and $(t,\omega)\in[0,T]\times\Omega$,
@@
$-\sum_{0<s\leq t}\left(g(X_{s}(\omega))-g(X_{s-}(\omega))-\nabla_{x}g(X_{s-}(\omega))\Delta X_{s}(\omega)\right)=M_{t}^{R}g(\omega)$
is an $(R,\mathbb{F})$-martingale as $R$ is a control rule in Definition 3.1. Moreover, we have, for $m\times R$-$\mathrm{a.s.}~{}(t,\omega)\in[0,T]\times\Omega$,
@@
We next study the existence of optimal relaxed controls. Consider the McKean-Vlasov SDE (1) and define the process by $Y_{t}:=X_{t}^{\alpha,\zeta}-\int_{0}^{t}G(s)\mathrm{d}\zeta_{s}$ for $t\in[0,T]$. Then, it follows from the dynamics (1) that
Motivated by this expression, we have the next result whose proof is omitted because the necessary part is obvious and the sufficiency can be shown by a classical martingale argument.
Lemma 3.5.
Let $R\in\mathcal{P}_{2}(\Omega;\widehat{\Omega})$. Then, $R\in\mathcal{R}_{ad}(\nu)$ if and only if there exists an $\mathbb{F}$-adapted process $Y=(Y_{t})_{t\in[0,T]}$ such that
  • (i)
    $Y$ has continuous trajectories and $R(X_{\cdot}=Y_{\cdot}+\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s})=1$;
  • (ii)
    $R\circ Y_{0}^{-1}=\nu$;
  • (iii)
  • (iv)
    for any $g\in C_{b}^{2}(\mathbb{R}^{n})$, $\tilde{M}_{t}^{R}g$ is a $(R,\mathbb{F})$-martingale, where
    $\tilde{M}_{t}^{R}g=g(Y_{t})-\int_{0}^{t}\int_{U}\tilde{\mathbb{G}}g(s,X_{s},R^{X}(s),u)(Y_{s})q_{s}(\mathrm{d}u)\mathrm{d}s$
    with the operator $\tilde{\mathbb{G}}g(t,x,\mu,u)(y):=b(t,x,\mu,u)\nabla_{x}g(y)+\frac{1}{2}\mathrm{tr}(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(y))$;
  • (v)
    for $i\in I$, $m\times R(A_{i}^{R})=T$.
Moreover, we have
Lemma 3.6.
The objective functional $\Gamma(R)$ defined by (5) is l.s.c. in $R\in\mathcal{R}_{\rm ad}(\nu)$.
Proof.
Then, for $R\in\mathcal{P}_{2}(\Omega;\widehat{\Omega})$, we get that $\Lambda(R,\cdot)$ is l.s.c. by virtue of Lemma 3.5 in Haussmann [20]. Furthermore, $\Lambda(R,\omega)$ is uniformly l.s.c. in $R\in\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ with respect to $\omega\in\Omega$ as $f$ is uniformly l.s.c. in $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ with respect to $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$. Thus, if $R_{\ell}\to R$ in $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ as $\ell\to\infty$ with $R_{\ell},R\in\mathcal{R}_{\rm ad}(\nu)$, then $R_{\ell}\to R$ in $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ as $\ell\to\infty$, and hence $R_{\ell}\to R$ as $\ell\to\infty$ with $R_{\ell},R\in\mathcal{R}_{\rm ad}(\nu)$. It then follows that
$\geq\liminf_{\ell\to\infty}\left\{\mathbb{E}^{R_{\ell}}[\Lambda(R_{\ell},\omega)]-\mathbb{E}^{R_{\ell}}[\Lambda(R,\omega)]\right\}+\liminf_{\ell\to\infty}\mathbb{E}^{R_{\ell}}[\Lambda(R,\omega)]$
$\geq\mathbb{E}^{R}[\Lambda(R,\omega)].$
The next result also holds.
Lemma 3.7.
The set $\{R\in\mathcal{P}_{2}(\Omega;\widehat{\Omega});~{}m\times R(A_{i}^{R})=T,~{}\forall i\in I\}$ is closed in $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$.
Proof.
For $R\in\mathcal{P}_{2}(\Omega;\Omega^{\prime})$ and $i\in I$, let us define a subset $\tilde{A}_{i}^{R}$ of $\Omega$ by
@@
We first show that $m\times R(A_{i}^{R})=T$ if and only if $R(\tilde{A}_{i}^{R})=1$.
Necessity. For any $\omega\in\Omega$, we define the following $\omega$-section that
@@
Then, it follows from Fubini’s Theorem that $T=m\times R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)$. This implies that there exists some $R$-null set $N\in\mathcal{F}$ such that $m(A_{i}^{R}(\omega))=T$ for all $\omega\in\Omega\backslash N$. For such $\omega$ and $0\leq t_{1}<t_{2}\leq T$, it holds that
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t=\int_{(t_{1},t_{2}]\cap A_{i}^{R}(\omega)}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t\geq 0.$
@@
Sufficiency. For $\omega\in\tilde{A}_{i}^{R}$, we have $m(A_{i}^{R}(\omega))=T$, and Fubini’s Theorem results in $m\times R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)=T$.
Now, we show that $\tilde{A}_{i}^{R}$ is closed for any $R\in\mathcal{P}_{2}(\Omega;\widehat{\Omega})$. Let $\omega^{\ell}=(\psi^{\ell},q^{\ell},\zeta^{\ell})\to\omega=(\psi,q,\zeta)$ in $\Omega$ as $\ell\to\infty$ and $\omega^{\ell}\in\tilde{A}_{i}^{R}$. Then, following the proof of Lemma 3.5 in Haussmann [20], we arrive at
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq\limsup_{\ell\to\infty}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi^{\ell}_{t}(\omega),R^{X}(t),u)q^{\ell}_{t}(\omega,\mathrm{d}u)\geq 0,$
where we have utilized the fact that $\phi^{i}$ is upper semi-continuous in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ imposed in Assumption 1-(A2). Hence, it holds that $\omega\in\tilde{A}_{i}^{R}$, which yields the closedness of $\tilde{A}_{i}^{R}$.
Let $R_{\ell}\to R$ in $\mathcal{P}_{2}(\Omega;\widehat{\Omega})$ as $\ell\to\infty$ with $R_{\ell}$ satisfying $R_{\ell}(\tilde{A}_{i}^{R_{\ell}})=1$ for all $\ell\geq 1$. Hence, $R_{\ell}(\tilde{A}_{i}^{R_{\ell}})-R_{\ell}(\tilde{A}_{i}^{R})=R_{\ell}(\tilde{A}_{i}^{R_{\ell}}\backslash\tilde{A}_{i}^{R})$. However, we have
$\tilde{A}_{i}^{R_{\ell}}\backslash\tilde{A}_{i}^{R}$$\subset\left\{\omega\in\Omega;~{}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi_{t}(\omega),R_{\ell}^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq 0,\right.$
$\qquad\left.\int_{U}\phi^{i}(t,\psi_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)<0,~{}\exists 0\leq t_{1}<t_{2}\leq T\right\}:=B_{\ell}.$
Note that $\phi^{i}$ is uniformly u.s.c. in $\mu$ with respect to $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$, which implies that
$\phi^{i}(t,\psi_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq\limsup_{\ell\to\infty}\phi^{i}(t,\psi_{t}(\omega),R_{\ell}^{X}(t),u)q_{t}(\omega,\mathrm{d}u),$
uniformly with respect to $(t,\omega,u)\in[0,T]\times\Omega\times\mathbb{R}^{l}$, and hence
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq\limsup_{\ell\to\infty}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R_{\ell}^{X}(t),u)q_{t}(\omega,\mathrm{d}u),$
uniformly with respect to $\omega\in\Omega$. Consequently, $B_{\ell}$ is an empty set whenever $\ell$ is sufficiently large. Thus, it holds that
@@
$=\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R_{\ell}}\backslash\tilde{A}_{i}^{R})+\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R})\leq R(\tilde{A}_{i}^{R}),$

### v2

@@ 5 Uniqueness and Stability of Solution to Constrained FBSDE
[…]
@@
$+\infty$ $>\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},\mathcal{L}^{R}(X_{t}),u)q_{t}(\mathrm{d}u)\mathrm{d}t\right]\geq\mathbb{E}^{R}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}^{R}(X_{t}),\alpha_{t})\mathrm{d}t\right]$
@@
This, together with Lemma 3.2, indicates that $\mathbb{E}^{R}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. On the other hand, using the construction of $\alpha$ again, we deduce that, for any $g\in C_{b}^{2}(\mathbb{R}^{n})$ and $(t,\omega)\in[0,T]\times\Omega$,
@@
$\qquad-\sum_{0<s\leq t}\left(g(X_{s}(\omega))-g(X_{s-}(\omega))-\nabla_{x}g(X_{s-}(\omega))\Delta X_{s}(\omega)\right)=M_{t}^{R}g(\omega)$
is an $(R,\mathbb{F})$-martingale as $R$ is a control rule in Definition 3.1. Moreover, we have, for $m\otimes R$-$\mathrm{a.s.}~(t,\omega)\in[0,T]\times\Omega$,
@@
Proof of Lemma 3.7.
[…]
@@
We first show that $m\otimes R(A_{i}^{R})=T$ iff $R(\tilde{A}_{i}^{R})=1$. Here, recall that $A_{i}^{R}\subset[0,T]\times\Omega$ is defined by (5).
Necessity. For any $\omega\in\Omega$, define the $\omega$-section of $A_{i}^{R}$ that
@@
Then, it follows from Fubini’s Theorem that $T=m\otimes R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)$. This implies that there exists some $R$-null set $N\in\mathcal{F}$ such that $m(A_{i}^{R}(\omega))=T$ for all $\N\omega\in\Omega\backslash$ N. For such $\omega$ and $0\leq t_{1}<t_{2}\leq T$, it holds that
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t$
$\qquad=\int_{(t_{1},t_{2}]\cap A_{i}^{R}(\omega)}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t\geq 0.$
@@
Sufficiency. For $\omega\in\tilde{A}_{i}^{R}$, we have $m(A_{i}^{R}(\omega))=T$, and Fubini’s Theorem results in $m\otimes R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)=T$.
[…]
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t\geq\limsup_{\ell\to\infty}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi^{\ell}_{t},R^{X}(t),u)q^{\ell}_{t}(\mathrm{d}u)\mathrm{d}t\geq 0,$
where we utilized the fact that $\phi^{i}$ is u.s.c. in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ imposed in Assumption 1-(A2). Hence, $\omega\in\tilde{A}_{i}^{R}$, which yields the closedness of $\tilde{A}_{i}^{R}$.
[…]
$\qquad\int_{t_{1}}^{t_{2}}\left.\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t<0,~\exists 0\leq t_{1}<t_{2}\leq T\right\}:=B_{\ell}.$
Note that $\phi^{i}$ is uniformly u.s.c. in $\mu$ w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$, which implies that
$\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)\geq\limsup_{\ell\to\infty}\phi^{i}(t,X_{t}(\omega),R_{\ell}^{X}(t),u)$
uniformly w.r.t. $(t,\omega,u)\in[0,T]\times\Omega\times\mathbb{R}^{l}$, and hence
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t\geq\limsup_{\ell\to\infty}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R_{\ell}^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t$
uniformly w.r.t. $\omega\in\Omega$. Hence, $B_{\ell}$ is an empty set whenever $\ell$ is sufficiently large. Thus, it holds that
@@
$=\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R_{\ell}}\backslash\tilde{A}_{i}^{R})+\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R})\leq R(\tilde{A}_{i}^{R}).$